: Compare simple and compound interest (compounded annually) on a principal of $15,000 at an annual rate of 3% for a 5-year period.
Answer #1

Principal =15000
Number of Periods =5
Rate =3%
Simple Interest =Principal *Rate *Time =15000*3%*5 =2250

Compound Interest =Principal*(1+r)^n-Principal =15000*(1+3%)^5-15000 =2389.11